OpenStudy (anonymous):

Use Gauss’s theorem to find the volume of the solid region bounded by the paraboloids z=9−x^2−y^2 and z=3x^2+3y^2−16.

OpenStudy (abb0t):

Gauss is theorem is also Divegence theorem, I believe. Which says: Let E be a simple solid region and S is te boundry surface of E with the positive orientation. Let f be a fector field whose components have cintinuous first order partial deriv. Then: \[\int\limits \int\limits_{S }^{ } F • dS = \int\limits \int\limits_{E }^{ } \int\limits dive F dV\]

OpenStudy (abb0t):

where F is a vector.

OpenStudy (abb0t):

So, I think you want to start by getting the divergence of the vector field firest.

OpenStudy (anonymous):

How do I find the vector field, if all im given is equations for parabaloids?